How do i solve A(0,0), B(8,6)
Answer by cleomenius(959)

I assume you are trying to fine the distance using the pythagorean theoem/
We will create a third point, (0,6)as point C.
(AC)^2 + (BC)^2 = (AB)^2
(8)^2 + (6)^2 = (AB)^2
64 + 36 = (AB)^2
100 = (AB)^2
10 = AB.
Cleomenius.